OpenStudy (anonymous):

michael bought a lawn mower at 30% off for 357, what was the original price of the lawn motor?

OpenStudy (lgbasallote):

510

OpenStudy (lgbasallote):

that's 357 divided by 70% by the way (70% because 30% was the one that was subtracted from the original price so the rate left is 70%)

OpenStudy (anonymous):

x-30% of X= 357 gives us 510 as original price.

OpenStudy (anonymous):

In a more simple equation form: x-0.3x = 357 0.7x = 357 x=357 / 0.7 x = 510